The volume control on a surround-sound amplifier is adjusted so the sound intensity level at the listening position increases from 23 to 61 dB. What is the ratio of the final sound intensity to the original sound intensity?
Answer #1

I = 10 log ( I / I0)

23 = 10  log ( I1 / I0)

61 =  10 log ( I2 / I0)

Subtracting we get

61 - 23 = 10 ( log ( I2/I1) )

3.8 = log (I2/I1)

I2/I1 = 10^(3.8)

I2 = 6309.57I1
Answer #2

B = 10(dB) log(\frac{I}{10^{-12}})

Given, B_{i} = 23dB , B_{f} = 62 dB

B_{i} = 10(dB) log(\frac{I_{i}}{10^{-12}}) = 23 dB

B_{f} = 10(dB) log(\frac{I_{f}}{10^{-12}}) = 61 dB

Subtracting both equation

10(dB) log(\frac{I_{f}}{10^{-12}}) - 10dB log(\frac{t_{i}}{10^{-12}}) = 61 dB - 23 dB

10(dB) log(\frac{I_{f}}{I_{i}})= 38 dB

log (\frac{I_{f}}{I_{i}}) = 3.8.

\frac{I_{f}}{I_{i}} = 10^{3.8}

\frac{I_{f}}{I_{i}} = 6309.57
Answer #3

I = 10 log ( I / I0)

23 = 10  log ( I1 / I0)

61 =  10 log ( I2 / I0)

Subtracting we get

61 - 23 = 10 ( log ( I2/I1) )

3.8 = log (I2/I1)

I2/I1 = 10^(3.8)

I2/I1 = 6309.57

I2 = 6309.57 I1